The specialty of an athlete on the women's track team is the pole vault. She has a mass of 46 kg and her approach speed is 8.6 m/s. When she is directly above the bar, her speed is 1.3m/s. Neglecting air resistance and any energy absorbed by the pole, determine the amount she has raised herself as she crosses the bar.
Answer #1

In this problem, the increase of her potential energy is equal to the decrease of her kinetic energy.

Initial KE = ½ * 46 * 8.6^2 = 1701.08J
Final KE = ½ * 46 * 1.3^2 = 38.87 J
∆ KE = 1701.08 – 38.87 = 1662.21 J

PE = 46 * 9.8 * h = 450.8 * h
450.8 * h = 1662.21
h = 1827.45 ÷ 568.4
This is approximately h= 3.687 meters